# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ III.3 Positive definite maps and Bochner’s theorem for finite inverse semigroup
where $A_{k}(\sigma,h):=\sum_{s\in D_{k}}\sigma(\lfloor s\rfloor)\otimes h_{s}^{\dagger}$ and $h:=\{h_{s}\}_{s\in D_{k}}$. Note that Eq. (79) is true for all the $\mathcal{D}$-classes $D_{k}$. Now, let us assume that $\widehat{\Phi}(\sigma)$ is positive semidefinite for all $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$. Since $A_{k}(\sigma,h)^{\dagger}A_{k}(\sigma,h)$ is positive semidefinite for any $D_{k}$, $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$ and any choice of $h\subseteq\mathbb{C}^{n}$, the right hand side of Eq. (79) is non-negative. Hence, $\sum_{s,t\in D_{k}}\left<\widetilde{\Phi}(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h_{t},h_{s}\right>\geq 0$ for all $\mathcal{D}$-classes $D_{k}$ implying that the map $\widetilde{\Phi}$ is positive definite. This completes the proof in one direction.
For the other direction, let us fix $\overline{\rho}_{k}^{0}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\rho^{0}_{k}\in\mathrm{IrrU}(G_{e_{k}})$ and choose $h^{0}$ as the subset consisting of the vectors $h^{0}_{s}$ of the form
@@
where $B\in M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})\otimes\mathbb{C}^{n}$ and $\operatorname{tr}_{d_{\overline{\rho}^{0}_{k}}}$ represents the trace over the space $M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})$. Writing $B$ as $B=\sum_{p}X_{p}\otimes x_{p}$, where $X_{p}\in M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})$ and $x_{p}\in\mathbb{C}^{n}$, we have
$(h_{s}^{0})^{\dagger}=\operatorname{tr}_{d_{\overline{\rho}^{0}_{k}}}\left((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right)^{\dagger}\otimes\mathbb{I})B\right)$ $=\sum_{c,d,p,k,l}\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}(((X_{p}))_{dc})_{lk}x_{p},$ (81)
where $\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}$ represents the complex conjugate of the $(l,k)$-th matrix element of the $(d,c)$-th block of the block matrix $\overline{\rho}^{0}_{k}(\lfloor s\rfloor)$. Now for $\sigma=\overline{\rho}_{k}$,
@@
$=\begin{cases}\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}^{0}}}B&\text{when}~\overline{\rho}_{k}=\overline{\rho}_{k}^{0},\\ 0&\text{when}~\overline{\rho}_{k}\neq\overline{\rho}_{k}^{0},\end{cases}$ (82)
where in the last equality, we have used Schur orthogonality relation (Eq. (71)). We put the above form of $A_{k}(\overline{\rho}_{k},h^{0})$ in Eq. (79) to arrive at
@@
Now if we assume that $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is positive definite then the left hand side of the above equation is non-negative. Since, $B$ and $\overline{\rho}_{k}^{0}$ are chosen arbitrarily, $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$. This completes the proof. $\Box$
The Stinespring dilation theorem for completely positive maps is a fundamental theorem in $C^{*}$ algebra. Stinespring dilation theorem also exists for positive definite maps from a group to a von Neumann algebra [19, 20]. Here we prove a Stinespring dilation theorem for positive definite maps on the contracted algebra $\mathbb{C}_{0}[S]$ of a finite inverse semigroup $S$.
Theorem 5.
(Stinespring Dilation) Let $S$ be a finite inverse semigroup and $\mathbb{C}_{0}[S]$ the the contracted algebra of $S$. Let $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ be a linear map. Then $\Phi$ is positive definite if and only if there exists a Hilbert Space $\mathcal{H}$, a bounded operator $V:\mathbb{C}^{n}\to\mathcal{H}$ and a *-homomorphism $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ such that for all $s\neq 0$
@@
where $\mathds{1}:=\sum_{e\in E(S)}\lfloor e\rfloor$ with $E(S)$ denoting the set of non-zero idempotents of $S$, is the identity of the contracted algebra $\mathbb{C}_{0}[S]$.
@@
It can be shown that the above defined inner product is well defined. With this inner product defined, the quotient space becomes an inner product space and as we are working on finite dimensions, it becomes a Hilbert space which we will be denoting by $\mathcal{H}:=\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}/\mathcal{N}$. Now define a linear map on $\mathcal{H}$ as
@@
where the element $\lfloor s\rfloor\cdot f\in\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}$ is defined by $\lfloor s\rfloor\cdot f:=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\lfloor s\rfloor\lfloor t\rfloor\otimes f_{t}$. The map $\pi$ extends linearly to a map $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$. One can simply verify that $\pi(\lfloor s\rfloor\lfloor t\rfloor)=\pi(\lfloor s\rfloor)\pi(\lfloor t\rfloor)$. Notice that due to the linearity of $\pi$, the relation $\pi(\lfloor s\rfloor\lfloor t\rfloor)=\pi(\lfloor s\rfloor)\pi(\lfloor t\rfloor)$ is consistent with the fact that $\lfloor s\rfloor\lfloor t\rfloor=0$ if $\mathrm{dom(s)}\neq\mathrm{ran}(t)$. Define another map $V:\mathbb{C}^{n}\to\mathcal{H}$ by
@@
where $\mathds{1}:=\sum_{e\in E(S)}\lfloor e\rfloor$ with $E(S)$ denoting the set of non-zero idempotents of $S$, is the identity of the contracted algebra $\mathbb{C}_{0}[S]$ (see [4]). Let $V^{*}:\mathcal{H}\to\mathbb{C}^{n}$ denotes the adjoint of $V$. Then we have
@@
Now using the definition of $V$ and the action of $V^{*}$ as provided above, we have:
@@
Also we have
@@
Now we prove that $\pi$ is a *-homomorphism:
@@
Hence $\pi(\lfloor u\rfloor)^{\dagger}=\pi(\lfloor u^{-1}\rfloor)$ for all $u\in S$, implying that $\pi$ is a *-homomorphism.
Now, we assume that a map linear $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ can be written as $\Phi(\lfloor s\rfloor)=V^{*}\pi(\lfloor s\rfloor)V$ for some Hilbert Space $\mathcal{H}$, a bounded operator $V:\mathbb{C}^{n}\to\mathcal{H}$ and a *-homomorphism $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$. Then we have
@@
III.4 Choi theorem on Completely positive maps and Bochner’s theorem on positive definite maps on finite dimensional matrix algebra
As discussed earlier, the the matrix algebra $M_{n}(\mathbb{C})$ is a contracted algebra of the inverse semigroup $S_{M}=\{e_{ij}\}_{i,j=1}^{n}\cup\{0\}$. So, from Theorem (5) it is clear that a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ is positive definite if and only if it completely positive. Now from Bochner’s Theorem (4) we have that a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ is positive definite if and only if its Fourier transform $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(M_{m}(\mathbb{C}))$. From the discussions of Proposition (B), the inverse semigroup $S_{M}$ has two $\mathcal{D}$-classes, namely $D_{0}=\{0\}$ and $D_{1}=\{e_{ij}\}_{i,j=1}^{n}$ with $D_{1}$ being the relevant one to us. Let us fix the idempotent $e_{11}$. Now the maximal subgroup at $e_{11}$ is $G_{e_{11}}:=\{e_{11}\}$. Also we have $\lfloor e_{ij}\rfloor=e_{ij}$, $\mathrm{ran}(e_{ij})=e_{ii}$, $\mathrm{dom}(e_{ij})=e_{jj}$, $p_{\mathrm{ran}(e_{ij})}=e_{i1}$, $p_{\mathrm{dom}(e_{ij})}=e_{j1}$. Now from the definition of the induced representation $\overline{\rho}_{k}$, we have
@@
Since $\rho_{1}$ is irreducible unitary representation of the singleton group $G_{e_{11}}=\{e_{11}\}$, we have $\rho_{1}(e_{11})=1$. Identifying $E_{e_{ii},e_{jj}}$ with $e_{ij}$, we have $\overline{\rho}_{1}(e_{ij})=e_{ij}$. From the definition of Fourier transform of $\Phi$ (see Eq. (41)),
@@
where in Eq. (96) we have used the fact that $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$. From Eq. (97) we see that $\widehat{\Phi}(\overline{\rho}_{1})$ is the Choi matrix of the linear map $\Phi$. Hence in the context of finite dimensional matrix algebra, Bochner’s theorem reduces to the Choi theorem on completely positive maps.
III.5 Relation between complete positivity of maps and positivity of their Fourier transforms
In the previous subsection, we have shown that the Bochner’s theorem on positive definite maps on matrix algebra reduces to the Choi theorem for completely positive(CP) maps. It should be noted that the Fourier transforms appeared in Bochner’s theorem are with respect to the induced representations arising from the irreducible unitary representations of the maximal subgroups of the finite inverse semigroup. So it interesting to know if the following statement is true: Let $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ be a completely positive map and let $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ be a representation. Then, is $\Phi$ completely positive if and only if $\widehat{\Phi}(\rho)$ is positive semidefinite? In the following theorem we provide a necessary and sufficient condition on the representation $\rho$ such that we have an affirmative answer to the question.

### v2

@@ 3.3 Positive Definite Maps and Bochner’s theorem for Finite Inverse Semigroups
where
$A_{k}(\sigma,h):=\sum_{s\in D_{k}}\sigma(\lfloor s\rfloor)\otimes h_{s}^{\dagger},$ (89)
with $h:=\{h_{s}\}_{s\in D_{k}}$. Observe that Eq. (88) is true for all $\mathcal{D}$-classes $D_{k}$. Now, assume that $\widehat{\Phi}(\sigma)$ is positive semidefinite for all $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$. Since $A_{k}(\sigma,h)^{\dagger}A_{k}(\sigma,h)$ is positive semidefinite for every choice of $D_{k}$, $\sigma$ and $h$, each term in the right hand side of Eq. (88) is non-negative. Consequently,
$\sum_{s,t\in D_{k}}\left<\widetilde{\Phi}(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h_{t},h_{s}\right>\geq 0$
for all $\mathcal{D}$-classes $D_{k}$. By Proposition 4, this implies that the map $\widetilde{\Phi}$ is positive definite. This completes the proof in one direction.
For the other direction, fix $\overline{\rho}_{k}^{0}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$ induced by $\rho^{0}_{k}\in\mathrm{IrrU}(G_{e_{k}})$ and choose $h^{0}$ as the subset consisting of the vectors $h^{0}_{s}$ of the form
@@
where $B\in M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})\otimes\mathbb{C}^{n*}$, $\operatorname{tr}_{d_{\overline{\rho}^{0}_{k}}}$ denotes the trace over the space $M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})$ and $\mathbb{C}^{n*}$ is the dual space of $\mathbb{C}^{n}$. Writing $B$ as a finite sum $B=\sum_{p}X_{p}\otimes x_{p}$, where $X_{p}\in M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})$ and $x_{p}\in\mathbb{C}^{n*}$, we have
$(h_{s}^{0})^{\dagger}$ $=\operatorname{tr}_{d_{\overline{\rho}^{0}_{k}}}\left((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right)^{\dagger}\otimes\mathbb{I})B\right)$ (91)
$=\sum_{c,d,p,k,l}\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}(((X_{p}))_{dc})_{lk}x_{p}.$ (92)
Here, $\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}$ denotes the complex conjugate of the $(l,k)$-th matrix element of the $(d,c)$-th block of the block matrix $\overline{\rho}^{0}_{k}(\lfloor s\rfloor)$. Now taking $\sigma=\overline{\rho}_{k}$, we have
@@
$=\begin{cases}\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}^{0}}}B&\text{when}~\overline{\rho}_{k}=\overline{\rho}_{k}^{0},\\ 0&\text{when}~\overline{\rho}_{k}\neq\overline{\rho}_{k}^{0}.\end{cases}$ (93)
The last equality follows from the Schur orthogonality relation for finite inverse semigroups, Eq. (80). Substituting this expression of $A_{k}(\overline{\rho}_{k},h^{0})$ into Eq. (88), we obtain
@@
Now assume that $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is positive definite, then the left hand side of the above equation is non-negative. Since $B$ and $\overline{\rho}_{k}^{0}$ were chosen arbitrarily, it follows that $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$. This completes the proof. $\Box$
The Stinespring dilation theorem for completely positive maps is a central result in $C^{*}$-algebra. The Stinespring dilation theorem also exists for positive definite maps from a group to a von Neumann algebra [45, 40]. Here we prove an analogous dilation theorem for positive definite maps on the contracted algebra $\mathbb{C}_{0}[S]$ of a finite inverse semigroup $S$.
Theorem 5 (Stinespring Dilation).
Let $S$ be a finite inverse semigroup and $\mathbb{C}_{0}[S]$ the the contracted algebra of $S$. Let $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ be a linear map. Then $\Phi$ is positive definite if and only if there exists a Hilbert Space $\mathcal{H}$, a bounded operator $V:\mathbb{C}^{n}\to\mathcal{H}$ and a *-homomorphism $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ such that for all $s\neq 0$
@@
where $\mathds{1}:=\sum_{e\in E(S)}\lfloor e\rfloor$ is the identity of the contracted algebra $\mathbb{C}_{0}[S]$, and $E(S)$ denotes the set of non-zero idempotents of $S$.
@@
It can be shown that the above defined inner product is well defined. With this inner product defined, the quotient space becomes an inner product space, and since we are working on finite dimensions, it is a Hilbert space which we denote by $\mathcal{H}:=\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}/\mathcal{N}$. Now, define a linear map on $\mathcal{H}$ as
@@
where the element $\lfloor s\rfloor\cdot f\in\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}$ is defined by
$\lfloor s\rfloor\cdot f:=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\lfloor s\rfloor\lfloor t\rfloor\otimes f_{t}.$ (100)
The map $\pi$ extends linearly to a map $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$. One can simply verify that
$\pi(\lfloor s\rfloor\lfloor t\rfloor)=\pi(\lfloor s\rfloor)\pi(\lfloor t\rfloor).$ (101)
Note that, due to the linearity of $\pi$, the relation $\pi(\lfloor s\rfloor\lfloor t\rfloor)=\pi(\lfloor s\rfloor)\pi(\lfloor t\rfloor)$ is consistent with the fact that $\lfloor s\rfloor\lfloor t\rfloor=0$ whenever $\mathrm{dom(s)}\neq\mathrm{ran}(t)$. Define another map $V:\mathbb{C}^{n}\to\mathcal{H}$ by
@@
where $\mathds{1}:=\sum_{e\in E(S)}\lfloor e\rfloor$ is the identity of the contracted algebra $\mathbb{C}_{0}[S]$ (see [14]). Let $V^{*}:\mathcal{H}\to\mathbb{C}^{n}$ denote the adjoint of $V$. Then we have
@@
Now, using the definition of $V$ and the action of $V^{*}$ as provided above, we obtain
@@
Thus
$\Phi(\lfloor s\rfloor)=V^{*}\pi(\lfloor s\rfloor)V$ (106)
Moreover,
@@
So,
$V^{*}V=\Phi\left(\mathds{1}\right).$ (108)
Now, we show that $\pi$ is a *-homomorphism. For $u\in S$ and $f],[h]\in\mathcal{H}$, we compute the following.
@@
Hence,
$\pi(\lfloor u\rfloor)^{\dagger}=\pi(\lfloor u^{-1}\rfloor)$ (110)
for all $u\in S$, which implies that $\pi$ is a *-homomorphism.
Conversely, suppose that a linear map $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ can be written as $\Phi(\lfloor s\rfloor)=V^{*}\pi(\lfloor s\rfloor)V$ for some Hilbert Space $\mathcal{H}$, a bounded operator $V:\mathbb{C}^{n}\to\mathcal{H}$ and a *-homomorphism $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$. Then for any subset $\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$, we have
@@
3.4 Choi’s Theorem as a Special Case of Bochner’s Theorem
As discussed earlier, the matrix algebra $M_{n}(\mathbb{C})$ is a contracted algebra of the inverse semigroup $S_{M}=\{e_{ij}\}_{i,j=1}^{n}\cup\{0\}$. Therefore, it follows from Theorem 5 that a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ is positive definite if and only if it completely positive. On the other hand, Bochner’s Theorem (4) for finite inverse semigroup implies that a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ is positive definite if and only if its Fourier transform $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(M_{m}(\mathbb{C}))$.
From Proposition B, the inverse semigroup $S_{M}$ has exactly two $\mathcal{D}$-classes, namely
$D_{0}=\{0\}\quad\text{and}\quad D_{1}=\{e_{ij}\}_{i,j=1}^{n},$
with $D_{1}$ being the only nontrivial class. Let us fix the idempotent $e_{11}$. Now, the maximal subgroup at $e_{11}$ is $G_{e_{11}}:=\{e_{11}\}$. Moreover, a straightforward computation shows
$\lfloor e_{ij}\rfloor=e_{ij},\quad\mathrm{ran}(e_{ij})=e_{ii},\quad\mathrm{dom}(e_{ij})=e_{jj},$
$p_{\mathrm{ran}(e_{ij})}=e_{i1},\quad and\quad p_{\mathrm{dom}(e_{ij})}=e_{j1}.$
Now from the definition of the induced representation $\overline{\rho}_{k}$ (see Theorem J), we obtain
@@ 3.4 Choi’s Theorem as a Special Case of Bochner’s Theorem
Since $\rho_{1}$ is an irreducible unitary representation of the singleton group $G_{e_{11}}=\{e_{11}\}$, we have $\rho_{1}(e_{11})=1$. Identifying $E_{e_{ii},e_{jj}}$ with $e_{ij}$, we have $\overline{\rho}_{1}(e_{ij})=e_{ij}$. Now, the definition of Fourier transform of $\Phi$ (see Eq. (50)) yields the following,
@@
where in Eq. (112) we have used the fact that $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$. From Eq. (113) we see that $\widehat{\Phi}(\overline{\rho}_{1})$ is precisely the Choi matrix of the linear map $\Phi$. Thus, in the context of finite dimensional matrix algebra, Bochner’s theorem reduces to the Choi theorem on completely positive maps.
3.5 Representations Characterizing Complete Positivity via Fourier Positivity
In Section 3.4, we showed that, for matrix algebras, Bochner’s theorem on positive definite maps reduces to the Choi theorem for completely positive (CP) maps. It should be noted that the Fourier transforms that appeared in Bochner’s theorem (Theorem 4) are taken with respect to the induced representations arising from the irreducible unitary representations of the maximal subgroups of the finite inverse semigroup. This naturally raises the question of what happens to the CP-positivity correspondence when a general representation of the matrix algebra is considered. More precisely, one may ask the following.
Let $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ be a completely positive map and let $\rho:M_{m}(\mathbb{C})\to M_{d_{\rho}}(\mathbb{C})$ be a representation. Then, is $\Phi$ completely positive if and only if its Fourier transform $\widehat{\Phi}(\rho)$ is positive semidefinite?
In the following theorem we provide a necessary and sufficient condition on the representation $\rho$ under which the answer to this question is affirmative.